An investor invested a total of ​$1,200 in two mutual funds. One fund earned a 8​% profit while the other earned a 6​% profit. If the​ investor's total profit was ​$90​, how much was invested in each mutual​ fund?
Found 2 solutions by ikleyn, greenestamps:
Answer by ikleyn(52780)

.

A TWIN problem was just solved today for you under this link

https://www.algebra.com/algebra/homework/Human-and-algebraic-language/Human-and-algebraic-language.faq.question.1199654.html

https://www.algebra.com/algebra/homework/Human-and-algebraic-language/Human-and-algebraic-language.faq.question.1199654.html


Read it attentively. Consider it as your TEMPLATE.

Solve the current problem and one million other similar problems in the same way.


Happy learning to you (!)



Answer by greenestamps(13200)



Assuming this post is from the same student who posted the other similar problem....

$1200 all invested at 6% would earn $72 interest; all at 8% would earn $96 interest; the actual interest was $90.

$90 is 18/24 = 3/4 of the way from $72 to $96, so 3/4 of the total was invested at the higher rate.

ANSWERS: 3/4 of $1200, or $900, was invested at 8%; the other $300 at 6%.

CHECK: .08(900)+.06(300) = 72+18 = 90